OpenStudy (nuccioreggie):

The temperature of a city changed by −24 degrees Celsius over a 6-week period. The temperature changed by the same amount each week. Part A: Write an expression to show the average change in the temperature of the city each week. (5 points) Part B: Simplify the expression and explain, using words, what your answer means. (5 points)

OpenStudy (nuccioreggie):

@paki

OpenStudy (nuccioreggie):

@mathmath333

OpenStudy (nuccioreggie):

@mathstudent55

OpenStudy (nuccioreggie):

@amistre64

OpenStudy (anonymous):

maybe -26/6= how much it changed pur week.. im not sure

OpenStudy (anonymous):

sry

OpenStudy (nuccioreggie):

@Michele_Laino

OpenStudy (michele_laino):

It is very simple, since, the average change, is given by the subsequent computation: -24/6 So, what is: -24/6=...?

OpenStudy (michele_laino):

@nuccioreggie

OpenStudy (nuccioreggie):

-4

OpenStudy (michele_laino):

ok!

OpenStudy (nuccioreggie):

Ok so thats the answer?

OpenStudy (michele_laino):

yes!

OpenStudy (nuccioreggie):

And can u please Simplify the expression and explain, using words, what your answer means.

OpenStudy (nuccioreggie):

But for part b?

OpenStudy (michele_laino):

for part B, you have to write that: "after a week, the change of temperature is equal to -4 degrees, so after a 6-weeks period, the total change of temperature is 4*(-6)=-24 degrees, as requested."

OpenStudy (nuccioreggie):

Tnak you maan so much can you just help me with one more

OpenStudy (michele_laino):

ok!

OpenStudy (nuccioreggie):

I ment thank you